OpenStudy (anonymous):

Solve the equation e^x+9e^-x/2=3 What is the solution in terms of natural logarithms? x= What is the decimal approximation for the solution? x=

OpenStudy (tkhunny):

It depends on what you mean. If you mean \(\dfrac{e^{x} + 9e^{-x}}{2} = 3\), then you should write that. You definitely have not.

OpenStudy (anonymous):

Yes, that is what I meant sorry...

OpenStudy (tkhunny):

Do you now why that is not what you have written?

OpenStudy (anonymous):

Refer to the attachment generated by Wolfram Alpha

OpenStudy (tkhunny):

Since I was a little crabby, I'll give you one for free... \(\dfrac{e^{x} + 9e^{-x}}{2} = 3\) 2 is never zero. Multiply by it. \(e^{x} + 9e^{-x} = 6\) \(e^{x}\) is never zero. Multiply by it \(e^{2x} + 9 = 6e^{x}\) Rearrange a little. \(e^{2x} - 6e^{x} + 9 = 0\) Recognize this as a quadratic-looking structure and simply factor it. \((e^{x} - 3)^{2} = 0\) Can you do the rest?

OpenStudy (anonymous):

So now simplify it?

OpenStudy (anonymous):

Is that the solution? How do I make it into a decimal?

OpenStudy (tkhunny):

You cannot simply that. We're solving aren't we? Do we have a value for 'x'? That would be a solution.

OpenStudy (tkhunny):

If you were solving a quadratic equation and encountered this, \((x-3)^{2} = 0\), what would you do?

OpenStudy (anonymous):

I am guessing some how put it into decimal form

OpenStudy (tkhunny):

You're saying that if you factor a quadratic equation, and get \((x-3)^{2} = 0\), you cannot solve it? x = What?

OpenStudy (anonymous):

Since I am having to put it into decimal terms and find the solution

OpenStudy (anonymous):

3

OpenStudy (tkhunny):

How did you get that?

OpenStudy (anonymous):

Solved the equation for x

OpenStudy (anonymous):

Is it wrong

OpenStudy (tkhunny):

Given \((x-3)^{2} = 0\), we conclude that \(x - 3 = 0\) or \(x = 3\). Is that what you mean by "solved the equation"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How would I put it into decimal form though?

OpenStudy (tkhunny):

Perfect. Now, do EXACTLY the same thing with \((e^{x} - 3)^{2} = 0\)

OpenStudy (anonymous):

decimal approximation**

OpenStudy (tkhunny):

We're not there, yet. Let's find the EXACT solution.

OpenStudy (tkhunny):

Remember that quadratic equation. Apply that EXACT technique to this equation.

OpenStudy (anonymous):

x=ln(3)

OpenStudy (anonymous):

or 1.9086

OpenStudy (tkhunny):

It helps a conversation if you go one step at a time. Very good. The REAL trick here was to recognize the Quadratic Form and treat it like that.

OpenStudy (anonymous):

Ok so which one is the exact solution and do I put 1.9086 for the decimal

OpenStudy (tkhunny):

What is not exact about ln(3)?